# A non-vanishing conjecture for cotangent bundles on elliptic surfaces

### v2

@@ 3.2. Local obstructions
Now let $A_{C}$ be an ample divisor on $C$ and $A_{E}$ be an ample divisor of degree one on $E$. Set $A=A_{C}\boxtimes A_{E}$ on $C\times E$. Choose a basis $s_{j,0},s_{j,1},\dots,s_{j,j-2},s_{j,j}$ of $H^{0}(E,\mathcal{O}_{E}(jA_{E}))$ such that each $s_{j,k}$ has vanishing order exactly $k$ at the origin $e\in E$. Consider a decomposition
$H^{0}(\Delta\times E,S^{m}\Omega_{\Delta\times E}^{1}\otimes\mathcal{O}_{\Delta\times E}(jA))=\bigoplus_{n}V_{m,n}$
@@
For $\omega\in H^{0}(\Delta\times E,S^{m}\Omega_{\Delta\times E}^{1}\otimes\mathcal{O}_{\Delta\times E}(jA))$, write $\omega=\sum_{n}\omega_{n}$ with $\omega_{n}\in V_{m,n}$.
@@
Remark that if $\omega\in\mathrm{im}(\Phi_{\Delta})$, then so is $\omega_{n}$ for each $n$ by the first part of [9, Lemma 6.11].
@@
In the chart $W_{01}$, we have
$s^{n-k}s_{j,k}ds^{\ell}dc_{0}^{m-\ell}=4^{-m}y_{0}^{\frac{n-3m}{4}}s_{0}^{n-k}(s_{0}dy_{0}+4y_{0}ds_{0})^{\ell}dy_{0}^{m-\ell}+O(y_{0}^{\frac{n-3m}{4}+1}),$
so one can apply the proof of [3, Section 3] to deduce that for any $\omega\in V_{m,n}$, $\mathrm{ord}_{\Theta_{0}}(\omega)>\frac{n-3m}{4}$ if and only if $y_{0}ds_{0}=c_{0}ds-sdc_{0}$ divides $\omega$. Hence for $\omega_{n}\in V_{m,n}\cap\mathrm{im}(\Phi_{\Delta})$, we have
@@
for some $\eta_{n}\in H^{0}(\Delta\times E,S^{\frac{m+n}{4}}\Omega_{\Delta\times E}^{1}\otimes\mathcal{O}_{\Delta\times E}((Nj-\frac{3m-n}{4})A))$. Now since $H^{0}(E,\mathcal{O}_{E}((Nj-\frac{3m-n}{4})A_{E}))=0$ if $Nj<\frac{3m-n}{4}$, the form $\omega_{n}$ is nonzero only when $n\geq 3m-4Nj$. ∎
@@ 3.3. The case of $II^{\ast}$, $III^{\ast}$ and $IV^{\ast}$
Consider $x\in Z_{III^{\ast}}$ and write $G_{x}=\langle g\rangle\simeq\mathbf{Z}/4\mathbf{Z}$. Again, choose a small disc $x\in\Delta\subset C$ as before. Replacing the origin of $E$ if necessary, the fixed points of $g$ are $p_{0}=(x,[0])$ and $p_{1}=(x,[\frac{1}{2}+\frac{1}{2}i])$; the fixed points of $g^{2}$ is $p_{2}=(x,[\frac{1}{2}])$. One can choose local coordinates $(s,c_{\nu})$ at $p_{\nu}$ such that $g:(s,c_{\nu})\mapsto(\eta_{4}s,\eta_{4}^{-1}c_{\nu})$ for $\nu=0,1$ and $g^{2}:(s,c_{2})\mapsto(-s,-c_{2})$. Hence the quotient $(\Delta\times E)/G_{x}$ has an $A_{3}$-singularity at $q_{\nu}=\pi(p_{\nu})$ if $\nu=0,1$ and an $A_{1}$-singularity at $q_{2}=\pi(p_{2})$. One can use the following result analogous to Lemma 3.2:
@@
In those charts, $\Theta_{\nu 1}$, $\Theta_{\nu 2}$ and $\Theta_{\nu 3}$ are defined by $y_{\nu}=s_{\nu 2}=0$, $t_{\nu 1}=s_{\nu 3}=0$ and $t_{\nu 2}=x_{\nu}=0$ respectively; $\Theta_{2}$ is defined by $y_{2}=x_{2}=0$; $\Theta$ is defined by $s_{\nu 1}=s_{2}=0$. Also, $\Theta$ does not meet $W_{\nu 2}$, $W_{\nu 3}$, $W_{\nu 4}$ and $W_{22}$.
@@
Assume on the contrary that $f^{\ast}\Omega_{B}^{1}(D)$ is not pseudoeffective. Then either $B\not\simeq\mathbb{P}^{1}$ or $\sum_{i=1}^{s}(1-\frac{1}{\nu_{i}})\geq 2$ holds. Thus $f^{\prime\ast}\Omega_{B}^{1}(D^{\prime})$ is not pseudoeffective where

### v3

@@ 3.2. Local obstructions
Now let $A_{C}$ be an ample divisor on $C$ and $A_{E}$ be an ample divisor of degree one on $E$. Set $A=A_{C}\boxtimes A_{E}$ on $C\times E$. Choose a basis $s_{j,0},s_{j,1},\dots,s_{j,j-2},s_{j,j}$ of $H^{0}(E,\mathcal{O}_{E}(jA_{E}))$ such that each $s_{j,k}$ has vanishing order exactly $k$ at the origin $e\in E$.
As before, consider $x\in Z_{III}$ and choose a disc $x\in\Delta\subset C$ small enough. Let $g$ be a generator of $G_{x}\simeq\mathbb{Z}/4\mathbb{Z}$, and $p_{0}\in\Delta\times E$ a fixed point of $G_{x}$. Then the singularity at $p_{0}$ is resolved by a $(-4)$-curve $\Theta_{0}$. Choose a local coordinate $(s,c_{0})$ at $p_{0}$ where $g$ acts by the multiplication by $\eta_{4}$. We have a decomposition
$H^{0}(\Delta\times E,S^{m}\Omega_{C\times E}^{1}\otimes\mathcal{O}_{C\times E}(jA))=\bigoplus_{n}V_{m,n}$
@@
For $\omega\in H^{0}(\Delta\times E,S^{m}\Omega_{C\times E}^{1}\otimes\mathcal{O}_{C\times E}(jA))$, write $\omega=\sum_{n}\omega_{n}$ with $\omega_{n}\in V_{m,n}$.
@@
Remark that if $\omega\in\mathrm{im}(\Phi_{\Delta})$, then so is $\omega_{n}$ for each $n$ by [3, Prop 3.3].
@@
Recall from the proof of Lemma 3.2 that there exists a coordinate chart $W_{01}$ on $S^{\prime}|_{\Delta}$ with local coordinate $(y_{0},s_{0})$ such that
$\begin{cases}s^{4}=y_{0}s_{0}^{4},\\ c_{0}^{4}=y_{0}\end{cases}$
and $\Theta_{0}$ is defined by $y_{0}=0$. In the chart $W_{01}$, we have
$s^{n-k}s_{j,k}ds^{\ell}dc_{0}^{m-\ell}=4^{-m}y_{0}^{\frac{n-3m}{4}}s_{0}^{n-k}(s_{0}dy_{0}+4y_{0}ds_{0})^{\ell}dy_{0}^{m-\ell}+O(y_{0}^{\frac{n-3m}{4}+1}).$
Thus by observing the leading terms with respect to $y_{0}$ and $ds_{0}$, one can deduce that for any $\omega\in V_{m,n}$, $\mathrm{ord}_{\Theta_{0}}(\omega)>\frac{n-3m}{4}$ if and only if $y_{0}ds_{0}=c_{0}ds-sdc_{0}$ divides $\omega$. Hence for $\omega_{n}\in V_{m,n}\cap\mathrm{im}(\Phi_{\Delta})$, we have
@@
for some $\eta_{n}\in H^{0}(\Delta\times E,S^{\frac{m+n}{4}}\Omega_{C\times E}^{1}\otimes\mathcal{O}_{C\times E}((Nj-\frac{3m-n}{4})A))$. Now since $H^{0}(E,\mathcal{O}_{E}((Nj-\frac{3m-n}{4})A_{E}))=0$ if $Nj<\frac{3m-n}{4}$, the form $\omega_{n}$ is nonzero only when $n\geq 3m-4Nj$. ∎
@@ 3.3. The case of $II^{\ast}$, $III^{\ast}$ and $IV^{\ast}$
Consider $x\in Z_{III^{\ast}}$ and write $G_{x}=\langle g\rangle\simeq\mathbf{Z}/4\mathbf{Z}$. Again, choose a small disc $x\in\Delta\subset C$ as before. Replacing the origin of $E$ if necessary, the fixed points of $g$ are $p_{0}=(x,[0])$ and $p_{1}=(x,[\frac{1}{2}+\frac{1}{2}i])$; the fixed points of $g^{2}$ is $p_{2}=(x,[\frac{1}{2}])$. One can choose local coordinates $(s,c_{\nu})$ at $p_{\nu}$ such that $g:(s,c_{\nu})\mapsto(\eta_{4}s,\eta_{4}^{-1}c_{\nu})$ for $\nu=0,1$ and $g^{2}:(s,c_{2})\mapsto(-s,-c_{2})$. Hence the quotient $(\Delta\times E)/G_{x}$ has an $A_{3}$-singularity at $q_{\nu}=\pi(p_{\nu})$ if $\nu=0,1$ and an $A_{1}$-singularity at $q_{2}=\pi(p_{2})$. Analogously to Lemma 3.2, we have:
@@
Figure 3.2. Coordinate charts on the $f^{\prime}$-fibre of type $III^{\ast}$.
In those charts, $\Theta_{\nu 1}$, $\Theta_{\nu 2}$ and $\Theta_{\nu 3}$ are defined by $y_{\nu}=s_{\nu 2}=0$, $t_{\nu 1}=s_{\nu 3}=0$ and $t_{\nu 2}=x_{\nu}=0$ respectively; $\Theta_{2}$ is defined by $y_{2}=x_{2}=0$; $\Theta$ is defined by $s_{\nu 1}=s_{2}=0$. Also, $\Theta$ does not meet $W_{\nu 2}$, $W_{\nu 3}$, $W_{\nu 4}$ and $W_{22}$. As before, we visualize the setting in Figure 3.2.
@@
Assume on the contrary that $f^{\ast}\Omega_{B}^{1}(D)$ is not pseudoeffective. Then both $B\simeq\mathbb{P}^{1}$ and $\sum_{i=1}^{s}(1-\frac{1}{\nu_{i}})<2$ holds. Thus $f^{\prime\ast}\Omega_{B}^{1}(D^{\prime})$ is not pseudoeffective where